OpenStudy (anonymous):

One more question, how do you solve 6 divided by 2/3= N?

OpenStudy (anonymous):

9

OpenStudy (anonymous):

Ok, Thanks but how?

OpenStudy (anonymous):

It's literally what you typed... 6 divided by (2/3). If you're not understanding the fraction division, remember to reciprocate the divisor, then multiply, like so: \[6/1 * 3/2 = 18/2 = 9\]

OpenStudy (anonymous):

Right! Thanks!

OpenStudy (anonymous):

6/1/2/3= (6*3)/(1*2)= 9

OpenStudy (anonymous):

Alexray, can you explain why N- 1/5 = 1/5?? Thanks!

OpenStudy (anonymous):

N would have to be double the quantity of 1/5 if subtracting 1/5 from it leaves 1/5. So N = 2/5

OpenStudy (anonymous):

Oh ok thanks,so out of these answer choses which would it be a. 3/20 b. 11/20 c.8/25 d. 6/15?

OpenStudy (anonymous):

choices O_O sorry im tired